# Uncertainty Quantification with the Empirical Neural Tangent Kernel

### v1

@@ Appendix B Related Works: Further Details and Discussions
Appendix C Hyper-parameter Tuning
@@
C.1 Regression
For regression, and with a sufficiently small learning rate, NUQLS samples from the distribution given in Remark 3.3. We can see that the variance of the predictions scales linearly with $\gamma^{2}$. Hence, we use the following framework to tune $\gamma$:
@@ C.1 Regression
    To obtain $\bm{\theta}_{s}$ in Algorithm 1, we initialise our parameters with a very small gamma, e.g. $\gamma=0.01$. This enables (stochastic) gradient descent to converge quickly with a small learning rate.
@@
    We compute $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},\mathcal{X}_{\text{val}})\}_{s=1}^{S}$, where $\mathcal{X}_{\text{val}}$ is the inputs from a validation set. As per Remark 3.3, for each point ${\bf x}\in\mathcal{X}_{\text{val}}$, we compute the mean prediction as $\mu(x)=\text{SampleMean}\left(\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},{\bf x})\}_{s=1}^{S}\right)$ and the variance as $\sigma^{2}_{\gamma}(x)=\text{SampleVariance}\left(\{\gamma\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},\mathcal{X}_{\text{val}})\}_{s=1}^{S}\right)$ (note the scaling by $\gamma$ for only the variance). We then use these values to compute the ECE across the validation dataset $\mathcal{D}_{\text{val}}$.
@@
    As coverage of a confidence interval scales linearly with the size of the given standard deviation, and our computed standard deviation scales linearly with $\gamma$, we find that the ECE is convex in $\gamma$. Hence, we employ the Ternary search method (see Algorithm 2) to find the value $\hat{\gamma}$ that minimizes this validation ECE.
@@
As can be seen in Table 1 and Table 6, this framework means that for regression our method is incredibly fast and computes well-calibrated variance values.
@@
  while $|l-r|\geq\delta$ and $i<$iter do
@@
C.2 Classification
For classification, with cross-entropy loss, there is no principled connection to a GP, nor a clean calibration method as in Section C.1. Instead, we propose some heuristics for how to choose the hyper-parameters of NUQLS. To accomplish this, we train a LeNet model on both MNIST and FashionMNIST, and run NUQLS on the trained model for a variety of learning rates, epochs and values of $\gamma$. We measure training loss of NUQLS, accuracy, ECE, AUC-ROC and OOD-AUC, as well as the training loss of the NN. We repeat this for $5$ random initialisations, and plot the average results in Figure 3.
Looking at Figure 3, we see that for loss values under the training loss of the original NN, performance is quite similar. Furthermore, for loss values over the training loss of the NN, a higher learning rate and number of epochs seems to equate to better performance. Finally, greater loss values, up to a threshold, seem to increase the OoD detection performance (measured by AUC-ROC and OOD-AUC). Greater loss values may equate to greater deviation in the row-space components of the ensemble solutions; this extra variance may help uncertainty quantification performance, though calibration for ID points seems to decrease (as measured by ECE).
We propose the following heuristic for hyper-parameter selection:
  • For a set learning rate and number of epochs, choose the largest $\gamma$ such that the NUQLS training loss is less than the NN training loss.
  • Parameter combinations such that the NUQLS training loss is moderately greater than the NN training loss may be better for uncertainty quantification purposes, thought the predictive calibration may be worse.
Figure 3: Comparison of NUQLS performance as a function of learning rate, epochs and standard deviation of initialisation $\gamma$. We train a LeNet model on (top) MNIST and (bottom) FashionMNIST, and compute a range of metrics. The results are averaged over $5$ random seeds. The blue horizontal line corresponds to the final training loss of the LeNet model.
Appendix D Further Experimental Results
D.1 Comparison to SLU

### v2

@@ B.2 SNGP
We also give mention to methods that seek to make DE more efficient, such as BatchEnsemble (BE) (Wen et al., 2020) and SnapshotEnsemble (Havasi et al., 2020). We note that these methods are yet to surpass DE as the SOTA in the literature. However,we compare NUQLS against BE in Figure˜6 and Table˜11, where we see that NUQLS shows superior performance, whilst remaining as scalable as BE. Further, BE requires modification to the structure of the network, and hence is not post-hoc.
B.4 Neural Tangent Kernel Methods
The Procedural-Noise-Correcting (PNC) method of Huang et al. (2023) employs the limiting NTK for infinite-width networks to characterize the noise in the optimization process of training a neural network. It uses this to provide a frequentist confidence interval around test predictions of the neural network, and as it arises from the frequentist framework, PNC recovers statistical guaranetees. In comparison, NUQLS does not rely on the limiting regime of the NTK, but rather employs the feature extraction of the empirical NTK to quantify uncertainty using a Bayesian framework. In Section˜H.1 we compare the performance of NUQLS vs. PNC. Unsurprisingly, given their constructions, we see that PNC outperforms NUQLS when the width of a network is extreme, while for finite-width networks that are fully trained, NUQLS outperforms PNC. This experiment illustrates that the methods are actually complementary, as they both excel in different settings.
Appendix C Uncertainty vs. Prediction
Here we discuss the use of common metrics to measure UQ ability in the BDL community. Generally, NLL, ECE and AUCROC (or the entropic version OODAUC) are used to measure the ability of a method to correctly quantify the uncertainty in a neural network. However, we argue that the use of these metrics for UQ is ill-advised, due to either a misalignment between the metric and the goal of UQ, or due to flawed measurements of uncertainty implicit in the metric. Specifically, we argue that variance over the softmax predictions is the uncertainty in the network, in the same way that the variance over probabilities is the uncertainty in a Dirichlet model. In contrast, NLL and AUCROC (OODAUC) are metrics for prediction, while ECE is a metric based on a flawed surrogate for uncertainty, calibration.
Prediction
While we will argue that uncertainty quantification must involve a computed variance term, and that common UQ metrics are either actually prediction metrics, or are based on a prediction of uncertainty, this discussion also begs the question: what about the purported benefits of BDL to prediction ability? As has been argued in Abe et al. (2022), it is often more economical to simply train a larger capacity model, than to use BDL to improve prediction quality. Hence, we restrict our main focus of NUQLS to its UQ ability. However, for completeness, we also evaluate the novel predictor of NUQLS; these results can be found in the Appendix˜G.
Uncertainty
Uncertainty can be divided into two categories: aleatoric uncertainty, or data uncertainty, and epistemic uncertainty, or model uncertainty. Aleatoric uncertainty arises from intrinsic randomness or noise in the data, and can often not be controlled. In contrast, epistemic uncertainty arises from having many potential models that may fit the data, and the uncertainty in not knowing which model is ‘correct’. As this arises from a lack of knowledge of the original data-mapping, uncertainty in this manner can often be reduced through increased training data, using a larger model, training for longer etc. There are numerous benefits to untangling the aleatoric and epistemic uncertainty (Mukhoti et al., 2021; Huseljic et al., 2021; Chan et al., 2024). We can use the framework of a Dirichlet model, the conjugate prior of a Categorical distribution, to understand how to evaluate these uncertainties in the context of $K$-class classification. Assume there is some subspace $\mathcal{M}$ of the parameter space $\bm{\Theta}\in^{p}$, such that our network ${\bf f}(.;\bm{\theta}^{*})$ attains ‘good’ test performance for any $\bm{\theta}^{*}\in\mathcal{M}$. If it is common across parameter values in $\mathcal{M}$ that ${\bf f}({\bf x};\bm{\theta}^{*})$ has high entropy for some $\bf x$}, i.e. ${\bf f}({\bf x};\bm{\theta}^{*})$ outputs a prediction close to the uniform distribution with high probability across $\bm{\theta}^{*}\in\mathcal{M}$, then we are sure that the outcome is unpredictable in $\bf x$}, i.e. that there is high aleatoric uncertainty. This equates to a Dirichlet model with parameters $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{K})$ where $\bm{\alpha}$ is uniform and of large magnitude. In contrast, assume that over all parameter values in $\mathcal{M}$, the softmax outputs of ${\bf f}({\bf x};\bm{\theta}^{*})$ for a given $\bf x$} have very large variance. This means that no model can agree on the ’correct’ data-mapping, and hence there is very large epistemic uncertainty. This is analogous to when the parameters $\bm{\alpha}$ of a Dirichlet model are small in magnitude. Our method, NUQLS, measures the variance over softmax outputs for high-performing parameters $\bm{\theta}^{*}\in\mathcal{M}$. Hence, we are measuring the epistemic uncertainty of the neural network.
NLL
Due to its probabilistic framework, NLL is often used as a metric for UQ ability. For $K$-class classification, up to a constant the NLL is just the cross-entropy loss
$\text{NLL}(\bm{\theta})=-\sum_{i=1}^{n}\sum_{k=1}^{K}y_{i,k}\log\mu_{i,k},$
where $y_{i,k}$ corresponds to the $k$-th element of the one-hot encoding of the $i$-th label, and $\mu_{i,k}=\text{softmax}({\bf f}({\bf x}_{i};\bm{\theta}))_{k}$. This loss is minimised as $\text{softmax}({\bf f}({\bf x}_{i};\bm{\theta}))_{k}\to\mathbbm{1}_{c_{i}}$, where $c_{i}$ is the correct label for input ${\bf x}_{i}$. While a lower NLL is generally considered to show better UQ ability, we can see that this metric is instead a smooth, probabilistic measurement of accuracy, where proximity of the predictor to the identity function is rewarded with lower NLL. This metric does not measure uncertainty quantification ability.
ECE
A common surrogate for uncertainty in neural networks is the predicted softmax probabilities. While neural networks generally output over-confident probabilities, many methods, such as temperature scaling and many Bayesian methods, attempt to create well-calibrated predictors. A well-calibrated output is one where the predicted softmax probabilities accurately convey the forecasted probability of seeing each of the $K$-classes. For example, of all test points where a network has a maximum predicted probability of $0.9$, we would hope that the network is $90\%$ accurate on these test points. Once a predictor is well-calibrated, the softmax probabilities are then seen to quantify uncertainty. However, there are issues with this:
  1. 1.
    Firstly, a point prediction from a network cannot be interpreted as the uncertainty in the network; uncertainty quantities must accompany a network prediction, for example with a variance term. At best, network confidence can be seen to predict the uncertainty.
  2. 2.
    Secondly, if softmax probability does in fact accurately predict the uncertainty, is this uncertainty due to data uncertainty, or model uncertainty? If a model accurately predicts that there is a $90\%$ chance that class $c$ is correct, does that uncertainty arise due to noise in the data, or some misspecification in the labeling process? Or is it due to the fact that the input is OoD, and thus the model is unsure of itself? Hence, we still require a variance term to ascertain what the epistemic uncertainty is.
  3. 3.
    Finally, calibration is only well-defined for ID test points. To see this, take a network trained on MNIST, and then tested on FashionMNIST. Any network prediction is meaningless on this test set, as the model has been trained for integer labels. The prediction that will minimize the calibration error will be the uniform prediction i.e. the prediction is $0.1,\dots,0.1]\in^{10}$ for all points in the OoD test set. However, the accuracy on OoD test points should be 0. Instead the predicted accuracy is $0.1$, which is what arises from randomly picking one class for each OoD test point. Therefore, the predicted probabilites cannot be directly interpreted as forecasted probability of accuracy for OoD points.
In practice, calibration is tested using ECE. This metric bins predictions, and then computes the average confidence in each bin. This is compared with the average predicted accuracy, and an $l_{2}$ norm is taken over the differences. As was noted in Nixon et al. (2019), there are several issues with ECE. For examples, it is reductive in a multi-class setting, where the bottom K1K-1 probabilities do not contribute meaningfully to the error, and it suffers from the inherent sharpness of neural networks, where the majority of maximum probabilities tend to be very close to $1$, and thus this confidence range is highly over-represented. So we struggle to even tell when a method is well-calibrated.
This issue with ECE can be summarised as follows: confidence is a flawed estimate of the total uncertainty; further, it is hard to tell when a method is well-calibrated. It is thus evident that it is much better to directly model the uncertainty in the model, by calculating the variance over the softmax predictions. Once equipped with this model uncertainty, entropy (or network confidence) can be better trusted to inform of the aleatoric uncertainty.
AUCROC
The final metric commonly used for UQ is AUCROC (OODAUC). This metric uses the top softmax prediction (entropy) as a predictor, to attempt to detect wether a point is ID or OoD. While this is a very useful task, it is still a metric of prediction, as only the mean predictor is used to compute this value. As we have argued, UQ requires a variance term. Miani et al. (2024a) in fact use the variance over logits to compute the AUCROC score, and report this score for their method SLU in comparison to other leading UQ methods. In Section˜H.2 we show that in this metric, we outperform SLU.
VMSP
To combat the issues of the previous metrics, we compare the performance of Bayesian methods using the variance of the maximum softmax predictor, or VMSP. We first provide an explicit definition of VMSP: for a Bayesian method, we generally have a mean predictor $\mu:\mathbb{R}^{d}\to\mathbb{R}^{c}$ and a covariance function $\Sigma:\mathbb{R}^{d}\to\mathbb{R}^{c\times c}$, for example the mean and covariance of the linearized ensemble in the case of NUQLS, that output in the probit space. To compute VMSP for a given test point $\mathbf{x}^{\star}$, we first find $\hat{c}=\text{argmax}_{k}~\mu(\mathbf{x}^{*})_{k}$, where $\mu(\mathbf{x}^{\star})_{k}$ denotes the $k$-th output of $\mu(\mathbf{x}^{\star})$. That is, we find the class that the Bayesian method predicts, given $\mathbf{x}^{\star}$. We then define $\text{VMSP}:=\Sigma(\mathbf{x}^{\star})_{\hat{c},\hat{c}}=\sigma^{2}(\mathbf{x}^{\star})_{\hat{c}}$, that is, the variance of this prediction.
We also introduce a pictorial method for evaluating the performance of Bayesian methods, using VMSP as the correct uncertainty measure. For a given dataset, we want a UQ method to provide low uncertainty for correctly predicted test points, and high uncertainty for incorrectly predicted or OOD test points. We then compare these distributions pictorially using a violin plot, and quantitatively using the median and skew values for the respective distributions. With the addition of a poorly-performing baseline model, we are able to easily compare the ability of UQ models to quantify uncertainty.
Appendix D Guarantees Against Mode Collapse
Given that all linearized models are initialized locally around the trained parameters, one may consider whether it is possible that all linear networks will converge to the same parameter solution, or mode of the linear loss. Fortunately, we are able to show that this will not occur, almost surely. To see this, we note that the solution to ˜3 for each linearized model is given by a unique row-space component (given our assumptions), plus a projection of the initialization ${\bf z}_{i}$ onto the null-space of the Jacobian. For mode-collapse, we would require the projection for two i.i.d initializations ${\bf z}_{1},{\bf z}_{2}\sim N(\mathbf{0},\gamma^{2}\mathbf{I})$ to be equal, or $(\mathbf{I}-{\mathbf{J}}_{\mathcal{X}}^{\dagger}{\mathbf{J}}_{\mathcal{X}})({\bf z}_{1}-{\bf z}_{2})=\mathbf{0}$, where ${\bf z}_{1}-{\bf z}_{2}\neq\mathbf{0}$ with probability one. Hence, ${\bf z}_{1}-{\bf z}_{2}={\mathbf{J}}_{\mathcal{X}}^{\dagger}{\mathbf{J}}_{\mathcal{X}}({\bf z}_{1}-{\bf z}_{2})$, and thus ${\bf z}_{1}-{\bf z}_{2}\in\text{Range}({\mathbf{J}}_{\mathcal{X}}^{T})$, i.e. ${\bf z}_{1},{\bf z}_{2}\in\text{Range}({\mathbf{J}}_{\mathcal{X}}^{T})$. However, $\text{Range}({\mathbf{J}}_{\mathcal{X}}^{T})$ is a low-dimensional subspace of $\mathbb{R}^{p}$, and as we are drawing i.i.d. from a $p$-dimensional normal distribution (i.e. not degenerate), the probability of this event occurring is 0. Hence, we do not need to be concerned with all models converging to the same behaviour.
Appendix E Computational Cost
[…]
Appendix F Hyper-parameter Tuning
@@
F.1 Regression
For regression, and with a sufficiently small learning rate, NUQLS samples from the distribution given in ˜3.3. We can see that the variance of the predictions scales linearly with $\gamma^{2}$. Hence, we use the following framework to tune $\gamma$:
@@ F.1 Regression
    To obtain $\bm{\theta}_{s}$ in Algorithm˜1, we initialise our parameters with a very small gamma, e.g. $\gamma=0.01$. This enables (stochastic) gradient descent to converge quickly with a small learning rate.
@@
    We compute $\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},\mathcal{X}_{\text{val}})\}_{s=1}^{S}$, where $\mathcal{X}_{\text{val}}$ is the inputs from a validation set. As per ˜3.3, for each point ${\bf x}\in\mathcal{X}_{\text{val}}$, we compute the mean prediction as $\mu(x)=\text{SampleMean}\left(\{\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},{\bf x})\}_{s=1}^{S}\right)$ and the variance as $\sigma^{2}_{\gamma}(x)=\text{SampleVariance}\left(\{\gamma\widetilde{{\bf f}}(\bm{\theta}^{\star}_{s},\mathcal{X}_{\text{val}})\}_{s=1}^{S}\right)$ (note the scaling by $\gamma$ for only the variance). We then use these values to compute the ECE across the validation dataset $\mathcal{D}_{\text{val}}$.
@@
    As coverage of a confidence interval scales linearly with the size of the given standard deviation, and our computed standard deviation scales linearly with $\gamma$, we find that the ECE is convex in $\gamma$. Hence, we employ the Ternary search method (see Algorithm˜2) to find the value $\hat{\gamma}$ that minimizes this validation ECE.
@@
As can be seen in Table˜1 and Table˜9, this framework means that for regression our method is incredibly fast and computes well-calibrated variance values.
@@
while $|l-r|\geq\delta$ and i<i<iter do
@@
F.2 Classification
For uncertainty quantification performance, as measured by VMSP in Figure˜3, we find that as long as $\gamma$ is small, training SGD for a small amount of epochs generally gives small training loss, and hence provides good performance. This means that our method can be computed quite quickly in larger data/model settings.
Appendix G Evaluation of Predictive Mean
While the main focus of NUQLS is the variance term, to compute the epistemic uncertainty of a NN, we would also like to demonstrate the predictive ability of our novel predictive mean, for completeness.
G.1 Image Classification - Predictive
Table 2: Image classification predictive performance, using LeNet5 on MNIST and FashionMNIST (FMNIST). Experiment was run 55 times with different random MAP initialisations to get standard deviation on metrics.
MAP $034\pm0.0$020.002 $\mathbf{0.990\pm}$$\mathbf{0.001}$ $008\pm0.0$010.001 $888\pm0.0$080.008 $886\pm0.0$080.008 $257$
NUQLS $035\pm0.0$020.002 $989\pm0.0$010.001 $\mathbf{0.003\pm}$$\mathbf{0.000}$ $\mathbf{0.930\pm}$$\mathbf{0.026}$ $\mathbf{0.928\pm}$$\mathbf{0.026}$ $106$
MNIST DE $034\pm0.0$040.004 $\mathbf{0.991\pm}$$\mathbf{0.000}$ $011\pm0.0$040.004 $\mathbf{0.932\pm}$$\mathbf{0.009}$ $\mathbf{0.928\pm}$$\mathbf{0.009}$ $2845$
MC-Dropout $044\pm0.0$020.002 $989\pm0.0$000.000 $7\pm0.010.01$ $873\pm0.0$320.032 $871\pm0.0$310.031 $533$
SWAG $\mathbf{0.029\pm}$$\mathbf{0.003}$ $\mathbf{0.991\pm}$$\mathbf{0.000}$ $\mathbf{0.004\pm}$$\mathbf{0.002}$ $902\pm0.0$080.008 $900\pm0.0$080.008 $489$
LLA* $034\pm0.0$020.002 $\mathbf{0.990\pm}$$\mathbf{0.001}$ $008\pm0.0$010.001 $888\pm0.0$080.008 $886\pm0.0$080.008 $45$
VaLLA $034\pm0.0$020.002 $\mathbf{0.990\pm}$$\mathbf{0.001}$ $008\pm0.0$010.001 $889\pm0.0$080.008 $886\pm0.0$080.008 $1583$
MAP $298\pm0.0$070.007 $0.891\pm$$0.3$ $\mathbf{0.006\pm}$$\mathbf{0.001}$ $840\pm0.0$220.022 $804\pm0.0$210.021 $158$
NUQLS $302\pm0.0$060.006 $891\pm0.0$020.002 $\mathbf{0.005\pm}$$\mathbf{0.002}$ $\mathbf{0.904\pm}$$\mathbf{0.007}$ $\mathbf{0.870\pm}$$\mathbf{0.006}$ $89$
FMNIST DE $288\pm0.0$020.002 $896\pm0.0$010.001 $013\pm0.0$010.001 $876\pm0.0$030.003 $836\pm0.0$030.003 $1587$
MC-Dropout $306\pm0.0$070.007 $892\pm0.0$030.003 $026\pm0.0$020.002 $856\pm0.0$210.021 $813\pm0.0$190.019 $291$
SWAG $\mathbf{0.283\pm}$$\mathbf{0.005}$ $\mathbf{0.899\pm}$$\mathbf{0.003}$ $018\pm0.0$020.002 $817\pm0.0$230.023 $783\pm0.0$220.022 $264$
LLA* $298\pm0.0$070.007 $891\pm0.0$030.003 $\mathbf{0.006\pm}$$\mathbf{0.001}$ $841\pm0.0$220.022 $805\pm0.0$210.021 $26$
VaLLA $298\pm0.0$070.007 $891\pm0.0$030.003 $007\pm0.0$010.001 $841\pm0.0$220.022 $805\pm0.0$210.021 $1583$
We compare NUQLS against the MAP NN, DE, MC-Dropout, SWAG, LLA* and VaLLA, on the MNIST and FashionMNIST datasets, using the LeNet5 network. We compare test cross-entropy (NLL), test accuracy (ACC), ECE, and the AUC-ROC measurement for maximum softmax probability (AUC-ROC) and entropy (OOD-AUC) as the detector. The last two metrics evaluate a methods ability to detect out-of distribution points. We display the results in Table˜2. We see that NUQLS performs the best in ECE, AUC-ROC and OOD-AUC, and is competitive in the other metrics, while having the second fastest wall-time.
Appendix H Further Experimental Results
H.1 Comparison to PNC using Confidence Intervals
We use this section to compare the difference in performance of NUQLS vs PNC. To test the difference between the two methods, we tested NUQLS on the confidence intervals problem from Huang et al. (2023). In this experiment, an MLP with a single hidden-layer is trained on $n$ datapoints from $U([0,0.2]^{d})$, where the target is $y=\sum_{i=1}^{d}sin(x_{i})+\epsilon$, and $\epsilon$ is a small noise term. The method is then asked to form a confidence interval around the prediction for ${\bf x}=(0.1,0.1,\dots,0.1)$. This setup is repeated several times, i.e. the network and training data are randomly initialized, and the coverage and average width of the confidence intervals are recorded, as well as the mean prediction. For more details, we refer the reader to Huang et al. (2023). When the width of the network is $32\times n$, and the network is only partially trained with $80$ epochs and a learning rate of $0.01$, NUQLS performs poorly compared to PNC, as can be seen in Table˜3. Note that bolded numbers indicate intervals that have reached or exceeded the expected coverage. Further, smaller width intervals are preferred.
Table 3: Evalaution of coverage (CR) and width (IW) of computed confidence intervals from toy problem, for a neural network with extreme width and partial training. A computed coverage exceeding the expected coverage is bolded. The mean prediction (MP) is also provided.
PNC NUQLS
$95\%$CI (CR/IW) $90\%$CI (CR/IW) MP $95\%$CI (CR/IW) $90\%$CI (CR/IW) MP
d=2,n=128d=2,~n=128 $\mathbf{0.98}/0.0437$ $\mathbf{0.95}/0.0323$ $0.1998$ $0.93/0.0357$ $\mathbf{0.92}/0.0299$ $0.2047$
d=4,n=256d=4,~n=256 $\mathbf{0.98}/0.0411$ $\mathbf{0.95}/0.0304$ $0.3991$ $0.92/0.0596$ $0.86/0.0500$ $0.4084$
Table 4: Evalaution of coverage (CR) and width (IW) of computed confidence intervals from toy problem, for a neural network with finite width and full training. A computed coverage exceeding the expected coverage is bolded. The mean prediction (MP) is also provided.
PNC NUQLS
$95\%$CI (CR/IW) $90\%$CI (CR/IW) MP $95\%$CI (CR/IW) $90\%$CI (CR/IW) MP
d=2,n=128d=2,~n=128 $0.8/0.136$ $0.72/0.0105$ $0.2022$ $\mathbf{0.99}/0.0135$ $\mathbf{0.96}/0.0114$ $0.2012$
d=4,n=256d=4,~n=256 $\mathbf{0.96}/0.0437$ $\mathbf{0.90}/0.0336$ $0.4045$ $\mathbf{0.97}/0.0313$ $\mathbf{0.97}/0.0313$ $0.4030$
d=8,n=512d=8,~n=512 $0.88/0.0740$ $0.88/0.0568$ $0.8078$ $\mathbf{1.00}/0.0667$ $\mathbf{0.98}/0.0559$ $0.8050$
d=16,n=128d=16,~n=128 $0.8/0.1443$ $0.8/0.1108$ $1.6265$ $\mathbf{1.00}/0.1350$ $\mathbf{0.98}/0.1133$ $1.6121$
We note that scaling the width of a network by $32\times n$ is rare in practice, and that such a wide network is difficult to train. If we instead form an MLP with width equal to the number of training points, and increase epochs to $100$ and learning rate to $0.5$, so that the network is properly trained, NUQLS far outperforms PNC, as can be seen in Table˜4. We note that for these experiments we tuned the $\gamma$ hyper-parameter for NUQLS on a small validation set. We conclude that NUQLS and PNC are in fact complementary methods. For infinite-width networks near initialization, PNC performs well while NUQLS struggles. Conversely, for finite-width networks trained to a minimum, NUQLS excels while PNC performs poorly. We believe the latter regime is more representative of practical scenarios, where NUQLS offers significantly better performance.
H.2 Comparison to SLU